# Single-site dissipation stabilizes a superconducting nonequilibrium steady state in a strongly correlated lattice

### v1

@@ V Disorder: Stability and Breakdown of Dissipative $\eta$-Pair ODLRO
To make the robustness statements operational (and easy to assess for finite systems), we focus on two complementary diagnostics in the steady state: the one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$ and, more stringently, the long-distance correlator plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. The latter directly probes the survival of ODLRO and is typically more sensitive to symmetry-breaking and leakage channels than local coherence. Throughout, we keep the rotated-jump channel fixed and introduce only  static  (quenched) perturbations in the Hamiltonian or in the jump parameters, so that the question is whether the NESS remains close to the clean +y+y-polarized fixed point on the disorder- averaged level and across system sizes accessible to exact evolution.
A useful organizing principle, consistent with the discussion in Secs. III and IV, is that two distinct structures underlie the preparation protocol, and disorder can attack either one: (i) the local holon–doublon (HD) closure $\mathcal{H}_{\mathrm{HD}}=\bigotimes_{i}\mathrm{span}\{|0\rangle_{i},|\uparrow\downarrow\rangle_{i}\}$, which is protected by the on-site $\eta_{i}^{2}$ algebra of the designed jump; and (ii) the global slow manifold $\mathcal{H}_{\Psi}$ (the doubled $\eta$-multiplet used for the triangular/gapped effective generator), which is the sector in which the coherent Hubbard contribution can be consistently eliminated to leading order. In the clean limit, these two mechanisms work together to produce an effectively classical pumping scenario: the rotated jump pins each local pseudospin along +y+y direction, and the Liouvillian gap protects the late-time attraction to the product dark state.
We consider quenched disorder added to the microscopic Hubbard model,
@@
while keeping the same rotated local jump operators $L_{j}=\sqrt{\gamma}\,\eta_{j}^{-}e^{i\frac{\pi}{2}\eta_{j}^{x}}$ on a chosen driven set. Numerically we quantify robustness by averaging over $N_{\mathrm{dis}}$ disorder realizations. For each disorder realization, we obtain the steady state either by evolving the Lindblad dynamics to the long-time limit or by directly solving for the zero eigenmode of the Liouvillian, and evaluate
$|\Phi_{m}|_{\mathrm{ss}}$ $\equiv\frac{1}{N}\,\mathbb{E}|\sum_{i=1}^{N}\mathrm{Tr}(\rho_{\mathrm{ss}}\eta_{i}^{+})|,$ (36)
$|C_{m}(r)|_{\mathrm{ss}}$ $\equiv\frac{1}{N}\,\mathbb{E}|\sum_{i=1}^{N}\mathrm{Tr}(\rho_{\mathrm{ss}}\eta_{i}^{+}\eta_{i+r}^{-})|,$ (37)
where $\mathbb{E}[\cdots]$ denotes the disorder average. In the ideal projected picture (Sec. III), the product dark state yields $|\langle\eta_{i}^{+}\rangle_{\mathrm{ss}}|=1/2$ and $|\langle\eta_{i}^{+}\eta_{j}^{-}\rangle_{\mathrm{ss}}|=1/4$ $(i\neq j)$. Consequently, the most stringent test of robustness is whether a nonzero plateau $|C_{m}(r\!\rightarrow\!N/2)|_{\mathrm{ss}}\approx 1/4$ persists in the long-distance limit.
@@ V.1 Disorder with negligible impact on the $\eta$-paired NESS
Figure 4: Robustness against interaction disorder $U_{i}=U+\delta U_{i}$ as a function of $W_{U}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$ and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Parameters: 1D chain with PBC, $t=\gamma$, $U=8/\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied to site 11. Both correlators remain close to their projected values over the explored range of $W_{U}/\gamma$, indicating that the $\eta$-paired NESS is weakly affected by interaction inhomogeneity.
@@
Figure 5: Robustness against dissipation-strength disorder $\gamma_{i}=\gamma+\delta\gamma_{i}$ as a function of $W_{\gamma}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$, and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Parameters: 1D chain with PBC, $t=\gamma$, $U=8/\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied to site 11. Both correlators remain close to their projected values over a broad range of $W_{\gamma}/\gamma$, indicating that dissipative inhomogeneity weakly renormalizes the projected generator $\mathcal{L}_{\mathrm{eff}}^{(0)}$. Error bars denote the standard error of the mean.
@@
Figure 6: Robustness against Zeeman-field disorder $B_{i}^{z}\,S_{i}^{z}+B_{i}^{x}\,S_{i}^{x}$ as a function of $W_{Z}/\gamma$. (a1) $|\Phi_{m}|_{\mathrm{ss}}$, and (b1) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Parameters: 1D chain with PBC, $t=\gamma$, $U=8/\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied to site 11. Both correlators remain close to the projected values over the explored range of $W_{Z}/\gamma$, indicating that the $\eta$-paired NESS is only weakly affected by Zeeman inhomogeneity.
@@
Figure 7: Robustness against bond-disorder $\delta t_{ij}$ as a function of $W_{t}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$, and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Parameters follow those in Fig. 6. In the perturbative window (large-$U$ separation and/or Zeno pumping), bond-disorder-induced excursions into the singlon sector remain short-lived virtual processes and can be eliminated by a controlled Schur-complement expansion, resulting primarily in renormalized relaxation pathways within the reduced holon–doublon manifold. Consistently, both correlators stay pinned close to the projected values over a broad range of $W_{t}/\gamma$. For stronger bond disorder, fluctuations increase and the long-distance plateau is gradually suppressed, indicating the onset of real singlon population and the breakdown of the projected description.
@@
A crucial subtlety, however, is that in the particle–hole symmetric Hubbard setting on a bipartite lattice, bond inhomogeneity of A–B hopping (without frustrating same-sublattice hopping) does not induce coherent dynamics within the $\eta$-multiplet: the projected Hamiltonian contribution vanishes, $P[-iH_{\mathrm{Hub}},\cdot]P=0$ on $\mathcal{H}_{\Psi}\otimes\mathcal{H}_{\Psi}^{\ast}$, and therefore the leading projected generator $\mathcal{L}_{\mathrm{eff}}^{(0)}$ retains the same dark-state structure (Appendix B). Consequently, bond disorder does not tilt the dissipative axis selection inside the reduced manifold in the way that, e.g., angle disorder does.
Bond disorder affects the dynamics instead through the off-manifold blocks in the Schur-complement elimination, $P\mathcal{L}Q$ and $Q\mathcal{L}P$, which encode leakage pathways into singlons. Increasing the spread of $t_{ij}$ enhances and spatially inhomogenizes these couplings, thereby strengthening higher-order corrections of the form $P\mathcal{L}Q\,(Q\mathcal{L}Q)^{-1}Q\mathcal{L}P+\cdots$. In the perturbative regime where such excursions remain short-lived (set here by the dissipative separation provided by the driven sites), this mainly reshuffles relaxation pathways and can broaden transient time scales without qualitatively changing the attractor. Consistent with this picture, both $|\Phi_{m}|_{\mathrm{ss}}$ and the long-distance plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ remain close to the projected benchmarks $1/2$ and $1/4$ over a broad range of $W_{t}$ (up to finite-size corrections); see Fig. 7.
This robustness is not unlimited. When bond disorder becomes sufficiently strong that hopping-induced excursions are no longer perturbative relative to the dissipative separation scale, real singlon population can accumulate and the projected description begins to break down. Numerically, this crossover shows up as a noticeable suppression of the ODLRO plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$, typically more pronounced than the reduction of the one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$, together with increased sample-to-sample fluctuations at large $W_{t}$, as seen in Fig. 7.
@@ V.2 Disorder that suppresses or destroys the $\eta$-pair ODLRO plateau
Figure 8: Response to on-site (chemical-potential) disorder $\mu_{i}=\mu+\delta\mu_{i}$ as a function of $W_{\mu}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$, and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Parameters: 1D chain with PBC, $t=\gamma$, $U=6/\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied on site 11. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. In contrast to symmetry-compatible perturbations, on-site inhomogeneity acts as an effective longitudinal field in the $\eta$ sector and competes with the dissipative phase locking along +y+y. Consistently, both correlators decrease rapidly with increasing $W_{\mu}/\gamma$, and the long-distance plateau is strongly suppressed at moderate disorder, indicating the breakdown of the projected-manifold picture and the loss of robust ODLRO.
@@
At the level of a single site, $H_{\mu}$ acts as a random longitudinal field in the local $\eta$-doublet, i.e., it detunes the holon and doublon components through $\eta_{i}^{z}$. By itself this remains diagonal in the local Fock basis and does not create singlons. The crucial point is that, in the full microscopic Hubbard model, the hopping $H_{t}$ connects holon/doublon configurations to intermediate singly occupied ones, and the energetic denominators controlling these virtual excursions are set by the local many-body splittings. Potential disorder reshuffles these splittings site by site; consequently, as $W_{\mu}$ increases it becomes increasingly likely to find resonant (or near-resonant) local configurations for which the singlon-assisted processes are no longer suppressed by large denominators. In other words, $H_{\mu}$ does not directly create singlons, but it substantially enhances the efficiency of singlon-mediated mixing generated by $H_{t}$ by reducing (and in rare regions, essentially removing) the protection afforded by the interaction gap and/or by the dissipative separation.
From the viewpoint of the effective-generator expansion (Schur-complement elimination), this is precisely the regime where the off-manifold resolvent becomes large: the operator $(Q\hat{\mathcal{L}}Q)^{-1}$ ceases to be parametrically small because $H_{\mu}$ can pull parts of the $Q$-sector spectrum toward zero, turning previously higher-order corrections into leading-order contributions. Once this occurs, the rotated dissipation can no longer be viewed as an approximately closed pumping problem within the HD (or $\eta$-paired) manifold, and the attractive fixed point is destabilized by real leakage into singlon configurations.
Numerically, this mechanism manifests as a pronounced suppression of both the steady-state one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$ and the long-distance plateau $|C_{m}(r\rightarrow N/2)|_{\mathrm{ss}}$ with increasing $W$, and destroying ODLRO; see Fig. 8.
@@
Figure 9: Response to a random transverse field acting directly in the holon–doublon manifold, $H_{\perp}=\sum_{i}(h_{i}^{x}\,\eta_{i}^{x}+h_{i}^{y}\,\eta_{i}^{y})$ with $h_{i}^{x,y}\in[-W_{\perp},W_{\perp}]$, as a function of $W_{\perp}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$, and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Parameters: 1D chain with PBC, $t=\gamma$, $U=6/\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied on site 11. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Unlike potential disorder, $H_{\perp}$ is already operative within $\mathcal{H}_{\mathrm{HD}}$: it coherently rotates the local $\eta$ pseudospin away from the dissipatively selected +y+y axis. Consistently, both correlators are strongly reduced with increasing $W_{\perp}/\gamma$, demonstrating direct competition with the dissipative phase locking and a rapid suppression of the ODLRO plateau.
@@
$h_{i}^{x},\;h_{i}^{y}$ $\in$ $[-W_{\perp},\,W_{\perp}],$ (51)
provides a qualitatively different (and more dangerous) perturbation than potential disorder. While potential disorder becomes destructive mainly by enhancing singlon-mediated leakage once combined with hopping, $H_{\perp}$ acts already inside $\mathcal{H}_{\mathrm{HD}}$: it coherently rotates the local $\eta$ pseudospin away from the dissipatively selected +y+y axis.
Numerically, this competition is immediately visible. As soon as $H_{\perp}\neq 0$, both $|\Phi_{m}|_{\mathrm{ss}}$ and the long-distance plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ depart from the projected benchmarks and decrease monotonically with increasing $W_{\perp}$ [Fig. 9]. In particular, the ODLRO plateau $C_{\infty}\simeq 1/4$ characteristic of the clean projected theory is lost already at the smallest accessible $W_{\perp}$, indicating that the $\eta$-paired NESS is not perturbatively stable against transverse fields.
The microscopic reason is simple: the clean protocol relies on dissipative axis selection, i.e., local jumps pin each site toward the +y+y dark state. A transverse Hamiltonian field continuously generates coherent precession within $\mathcal{H}_{\mathrm{HD}}$, so the dark condition is no longer dynamically preserved and the steady state is pushed away from the +y+y product fixed point. The resulting NESS becomes a mixed state set by the competition between coherent rotation and dissipative repumping, with strongly reduced long-distance $\eta$ coherence.
@@
Figure 10: Response to angle disorder in the rotated dissipation, $\theta_{j}=\pi/2+\delta\theta_{j}$ with $\delta\theta_{j}\in[-W_{\theta},W_{\theta}]$, as a function of $W_{\theta}/\pi$. (a) $|\Phi_{m}|_{\mathrm{ss}}$ and, (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Parameters: 1D chain with PBC, $t=\gamma$, $U=6/\gamma$, nominal rotation $\theta=\pi/2$, $N=5$, and dissipation applied on site 11. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Angle disorder tilts the locally selected dissipative axis: the clean +y+y dark state is no longer exactly annihilated for $\delta\theta_{j}\neq 0$, so the attractive fixed point is shifted. Consistent with this, the steady-state correlators remain close to the projected values at weak disorder (small $W_{\theta}/\pi$), but are progressively suppressed as $W_{\theta}/\pi$ increases, with the long-distance correlator $|C_{m}|_{\mathrm{ss}}$ being more sensitive than the one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$.
Imperfect feedback rotation,
$L_{j}$ $=$ $\sqrt{\gamma}\,\eta_{j}^{-}e^{i\theta_{j}\eta_{j}^{x}},\qquad$ (52)
$\theta_{j}$ $=$ $\frac{\pi}{2}+\delta\theta_{j},\text{ }\delta\theta_{j}\in[-W_{\theta},\,W_{\theta}]$ (53)
is a direct perturbation of the pumping mechanism itself: it rotates the jump away from the ideal $\pi/2$ frame and therefore changes the locally selected dissipative axis. Equivalently, the clean dark state $|\eta_{j}^{y}=+1/2\rangle$ is no longer annihilated when $\delta\theta_{j}\neq 0$. As a consequence, even weak angle disorder immediately shifts the attractive fixed point and the NESS ceases to be the +y+y product state. In numerics this shows up as a rapid departure of both $|\Phi_{m}|_{\mathrm{ss}}$ and the long-distance plateau $|C_{m}(r\!\rightarrow\!N/2)|_{\mathrm{ss}}$ from their projected values as $W_{\theta}$ is increased; see Fig. 10.

### v2

@@ V Disorder: Stability and Breakdown of Dissipative $\eta$-Pair ODLRO
To quantify robustness in a way that is transparent for finite systems, we track two steady-state diagnostics: the one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$ and, more stringently, the long-distance plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. The plateau is a direct test of ODLRO and is typically more sensitive than local coherence to symmetry breaking and to leakage out of the targeted manifold. In what follows, we keep the rotated-jump channel fixed and introduce only static (quenched) perturbations either in the Hamiltonian or in the jump parameters. The question is then whether, after disorder averaging, the NESS remains close to the clean +y+y-polarized fixed point for the system sizes accessible to exact time evolution.
A useful organizing principle, consistent with Secs. III and IV, is that the preparation protocol rests on two distinct structures, and disorder may compromise either one: (i) the local holon–doublon (HD) closure $\mathcal{H}_{\mathrm{HD}}=\bigotimes_{i}\mathrm{span}\{|0\rangle_{i},|\uparrow\downarrow\rangle_{i}\}$, which is enforced by the local $\eta$-pseudospin selection rules of the designed jump; and (ii) the global slow manifold $\mathcal{H}_{\Psi}$ (the doubled $\eta$-multiplet supporting the triangular and gapped effective generator), within which the coherent Hubbard contribution can be eliminated consistently to leading order. In the clean limit these ingredients work in concert: the rotated jump pins each local pseudospin close to the +y+y direction, while the dissipative gap controls the late-time relaxation into the (product-like) dark attractor.
We now add quenched disorder to the microscopic Hubbard model,
@@
while keeping the same rotated local jump operators $L_{j}=\sqrt{\gamma}\,\eta_{j}^{-}e^{i\frac{\pi}{2}\eta_{j}^{x}}$ on a chosen driven set. Numerically, we assess robustness by averaging over $N_{\mathrm{dis}}$ disorder realizations. For each realization, we obtain the steady state either by evolving the Lindblad dynamics to long times or by directly solving for the zero-eigenvalue mode of the Liouvillian, and we evaluate
$|\Phi_{m}|_{\mathrm{ss}}$ $\equiv\frac{1}{N}\,\mathbb{E}\Bigl|\sum_{i=1}^{N}\mathrm{Tr}\!\left(\rho_{\mathrm{ss}}\,\eta_{i}^{+}\right)\Bigr|,$ (35)
$|C_{m}(r)|_{\mathrm{ss}}$ $\equiv\frac{1}{N}\,\mathbb{E}\Bigl|\sum_{i=1}^{N}\mathrm{Tr}\!\left(\rho_{\mathrm{ss}}\,\eta_{i}^{+}\eta_{i+r}^{-}\right)\Bigr|,$ (36)
where $\mathbb{E}[\cdots]$ denotes the disorder average. In the ideal projected picture of Sec. III, the product dark state yields $|\langle\eta_{i}^{+}\rangle_{\mathrm{ss}}|=1/2$ and $|\langle\eta_{i}^{+}\eta_{j}^{-}\rangle_{\mathrm{ss}}|=1/4$ for $i\neq j$. Accordingly, the most stringent robustness criterion is the survival of a nonzero long-distance plateau: whether $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ remains close to $1/4$ as $r$ approaches the maximal separation.
@@ V.1 Disorder with negligible impact on the $\eta$-paired NESS
Figure 4: Robustness against interaction disorder $U_{i}=U+\delta U_{i}$ as a function of $W_{U}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$ and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Parameters: 1D chain with PBC, $t=\gamma$, $U=8\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied to site 11. Both correlators remain close to their projected values over the explored range of $W_{U}/\gamma$, indicating that the $\eta$-paired NESS is weakly affected by interaction inhomogeneity.
@@
Figure 5: Robustness against dissipation-strength disorder $\gamma_{i}=\gamma+\delta\gamma_{i}$ as a function of $W_{\gamma}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$, and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Parameters: 1D chain with PBC, $t=\gamma$, $U=8\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied to site 11. Both correlators remain close to their projected values over a broad range of $W_{\gamma}/\gamma$, indicating that dissipative inhomogeneity weakly renormalizes the projected generator $\mathcal{L}_{\mathrm{eff}}^{(0)}$. Error bars denote the standard error of the mean.
@@
Figure 6: Robustness against Zeeman-field disorder $B_{i}^{z}\,S_{i}^{z}+B_{i}^{x}\,S_{i}^{x}$ as a function of $W_{Z}/\gamma$. (a1) $|\Phi_{m}|_{\mathrm{ss}}$, and (b1) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Parameters: 1D chain with PBC, $t=\gamma$, $U=8\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied to site 11. Both correlators remain close to the projected values over the explored range of $W_{Z}/\gamma$, indicating that the $\eta$-paired NESS is only weakly affected by Zeeman inhomogeneity.
@@
Figure 7: Robustness against bond disorder $\delta t_{ij}$ as a function of $W_{t}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$ and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Points show disorder averages over 100 realizations, and dashed lines mark the projected benchmarks $1/2$ and $1/4$. Parameters are the same as in Fig. 6, but with no Zeeman field. In the perturbative window set by the large-$U$ separation and/or Zeno pumping, bond-disorder-induced excursions into the singlon sector remain short-lived and can be treated as virtual processes in a controlled Schur-complement expansion. Their main effect is therefore to renormalize relaxation pathways within the reduced holon–doublon manifold. Accordingly, both correlators stay close to the projected values over a wide range of $W_{t}/\gamma$. For stronger bond disorder, sample-to-sample fluctuations grow and the long-distance plateau is gradually suppressed, signaling the buildup of real singlon population and the breakdown of the projected description.
@@
A crucial subtlety is that, for the particle–hole symmetric Hubbard model on a bipartite lattice, inhomogeneous A–B hopping, as long as it does not introduce frustrating same-sublattice terms, still does not generate coherent dynamics within the $\eta$-multiplet. In the reduced description this appears as a vanishing projected Hamiltonian contribution, $P[-iH_{\mathrm{Hub}},\cdot]P=0$ on $\mathcal{H}_{\Psi}\otimes\mathcal{H}_{\Psi}^{\ast}$, so the leading projected generator $\mathcal{L}_{\mathrm{eff}}^{(0)}$ keeps the same dark-state structure (Appendix B). As a result, bond disorder does not tilt the dissipative axis selected inside the reduced manifold, in contrast to angle disorder.
Bond disorder mainly influences the dynamics through the off-manifold couplings that connect the target subspace to singlon sectors, i.e., the $P\mathcal{L}Q$ and $Q\mathcal{L}P$ blocks in the Schur-complement elimination. A broader spread of $t_{ij}$ strengthens and spatially inhomogenizes these couplings, which in turn enhances higher-order corrections such as $P\mathcal{L}Q\,(Q\mathcal{L}Q)^{-1}Q\mathcal{L}P+\cdots$. In the regime where these excursions remain perturbative and short-lived (here controlled by the dissipative separation induced by the driven sites), the main effect is to reroute relaxation and broaden transient time scales, while leaving the steady state essentially unchanged. Accordingly, both $|\Phi_{m}|_{\mathrm{ss}}$ and the long-distance plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ stay close to the projected benchmarks $1/2$ and $1/4$ over a wide range of $W_{t}$ (up to finite-size effects); see Fig. 7.
This robustness has a clear limit. Once bond disorder is strong enough that hopping-induced excursions are no longer perturbative on the dissipative separation scale, singlon population can build up and the projected description ceases to apply. Numerically, this crossover appears as a visible suppression of the ODLRO plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$—typically stronger than the reduction of the one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$—together with larger sample-to-sample fluctuations at large $W_{t}$, as shown in Fig. 7.
@@ V.2 Disorder that suppresses or destroys the $\eta$-pair ODLRO plateau
Figure 8: Response to on-site (chemical-potential) disorder $\mu_{i}=\mu+\delta\mu_{i}$ as a function of $W_{\mu}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$, and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Parameters: 1D chain with PBC, $t=\gamma$, $U=6\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied on site 11. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. In contrast to symmetry-compatible perturbations, on-site inhomogeneity acts as an effective longitudinal field in the $\eta$ sector and competes with the dissipative phase locking along +y+y. Consistently, both correlators decrease rapidly with increasing $W_{\mu}/\gamma$, and the long-distance plateau is strongly suppressed at moderate disorder, indicating the breakdown of the projected-manifold picture and the loss of robust ODLRO.
@@
At the single-site level, $H_{\mu}$ acts as a random longitudinal field within the local $\eta$ doublet and detunes the holon and doublon components through $\eta_{i}^{z}$. This term is diagonal in the local Fock basis, so by itself it cannot populate the singlon sector. Its impact arises in the full Hubbard model, where hopping $H_{t}$ couples the holon–doublon manifold to intermediate singly occupied states via virtual processes. The corresponding energy denominators are set by local many-body splittings, which are reshuffled by the site-dependent potential offsets. As the disorder strength $W_{\mu}$ increases, near-resonant configurations become more frequent and the singlon-assisted processes are no longer suppressed by large denominators. As a result, $H_{\mu}$ does not create singlons directly, but it strongly enhances the singlon-mediated mixing generated by $H_{t}$ by weakening, and in rare regions effectively removing, the protection provided by the interaction gap and by the dissipative separation.
From the viewpoint of the effective-generator expansion based on Schur-complement elimination, this is the regime where the off-manifold resolvent becomes large. The operator $(Q\hat{\mathcal{L}}Q)^{-1}$ is no longer parametrically small, because $H_{\mu}$ can shift parts of the $Q$-sector spectrum toward zero and thereby promote terms that were higher order in the expansion to leading contributions. Once this happens, the rotated dissipation can no longer be viewed as an approximately closed pumping problem within the holon–doublon manifold or the $\eta$-paired sector. The attractive fixed point is then destabilized by real leakage into singlon configurations.
Numerically, this crossover shows up as a clear suppression of both the steady-state one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$ and the long-distance plateau $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$ as $W_{\mu}$ is increased. In this regime the ODLRO plateau collapses, as illustrated in Fig. 8.
@@
Figure 9: Response to a random transverse field acting directly in the holon–doublon manifold, $H_{\perp}=\sum_{i}(h_{i}^{x}\,\eta_{i}^{x}+h_{i}^{y}\,\eta_{i}^{y})$ with $h_{i}^{x,y}\in[-W_{\perp},W_{\perp}]$, as a function of $W_{\perp}/\gamma$. (a) $|\Phi_{m}|_{\mathrm{ss}}$, and (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Parameters: 1D chain with PBC, $t=\gamma$, $U=6\gamma$, $\theta=\pi/2$, $N=5$, and dissipation applied on site 11. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Unlike potential disorder, $H_{\perp}$ is already operative within $\mathcal{H}_{\mathrm{HD}}$: it coherently rotates the local $\eta$ pseudospin away from the dissipatively selected +y+y axis. Consistently, both correlators are strongly reduced with increasing $W_{\perp}/\gamma$, demonstrating direct competition with the dissipative phase locking and a rapid suppression of the ODLRO plateau.
@@
$h_{i}^{x,y}$ $\in$ $[-W_{\perp},W_{\perp}],$ (50)
represents a qualitatively different and more dangerous perturbation than potential disorder. While potential disorder becomes destructive primarily by enhancing singlon-mediated leakage through hopping, $H_{\perp}$ operates directly within $\mathcal{H}_{\mathrm{HD}}$ by coherently rotating the local $\eta$-pseudospin away from the dissipatively selected +y+y axis.
Numerically, this competition is immediately evident. As soon as $H_{\perp}\neq 0$, both the steady-state one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$ and the long-distance plateau $|C_{m}(r\to N/2)|_{\mathrm{ss}}$ depart from the projected benchmarks, decreasing monotonically with increasing $W_{\perp}$, as shown in Fig. 9. Notably, the ODLRO plateau $C_{\infty}\simeq 1/4$ characteristic of the ideal projected theory is compromised even at small $W_{\perp}$, indicating that the $\eta$-paired NESS is not perturbatively stable against transverse fields.
The microscopic origin of this sensitivity is straightforward. The clean protocol relies on dissipative axis selection, where local jumps pin each site toward the +y+y dark state. A transverse Hamiltonian field continuously generates coherent precession within $\mathcal{H}_{\mathrm{HD}}$, which prevents the dark condition from being dynamically preserved and forces the steady state away from the +y+y product fixed point. The resulting NESS becomes a mixed state determined by the competition between coherent rotation and dissipative repumping, leading to a substantial reduction in long-distance $\eta$-coherence.
@@
Figure 10: Response to angle disorder in the rotated dissipation, $\theta_{j}=\pi/2+\delta\theta_{j}$ with $\delta\theta_{j}\in[-W_{\theta},W_{\theta}]$, as a function of $W_{\theta}/\pi$. (a) $|\Phi_{m}|_{\mathrm{ss}}$ and, (b) $|C_{m}(r\!\to\!N/2)|_{\mathrm{ss}}$. Parameters: 1D chain with PBC, $t=\gamma$, $U=6\gamma$, nominal rotation $\theta=\pi/2$, $N=5$, and dissipation applied on site 11. Points show disorder averages over 100100 realizations; dashed lines mark the projected benchmarks $1/2$ and $1/4$. Angle disorder tilts the locally selected dissipative axis: the clean +y+y dark state is no longer exactly annihilated for $\delta\theta_{j}\neq 0$, so the attractive fixed point is shifted. Consistent with this, the steady-state correlators remain close to the projected values at weak disorder (small $W_{\theta}/\pi$), but are progressively suppressed as $W_{\theta}/\pi$ increases, with the long-distance correlator $|C_{m}|_{\mathrm{ss}}$ being more sensitive than the one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$.
Imperfect feedback rotation, described by the modified jump operators
$L_{j}$ $=$ $\sqrt{\gamma}\,\eta_{j}^{-}e^{i\theta_{j}\eta_{j}^{x}},$ (51)
$\theta_{j}$ $=$ $\frac{\pi}{2}+\delta\theta_{j},\quad\delta\theta_{j}\in[-W_{\theta},W_{\theta}],$ (52)
represents a direct perturbation of the pumping mechanism itself. This disorder rotates the jump operator away from the ideal $\pi/2$ frame, thereby shifting the locally selected dissipative axis. Consequently, the ideal dark state $|\eta_{j}^{y}=+1/2\rangle$ is no longer annihilated when $\delta\theta_{j}\neq 0$.
Even weak angle disorder immediately displaces the attractive fixed point, causing the NESS to depart from the pure +y+y product state. In numerical simulations, this mechanism manifests as a rapid decrease in both the steady-state one-point amplitude $|\Phi_{m}|_{\mathrm{ss}}$ and the long-distance plateau $|C_{m}(r\rightarrow N/2)|_{\mathrm{ss}}$ as $W_{\theta}$ increases, as illustrated in Fig. 10.